What principal should be invested today to have $3000 after two years, if the interest is paid at 6% per year, compounded quarterly? Show your work using a formula
Answer by Theo(13342)

f = p * (1+r) ^ n
f is th future value
p is the presnt value
r is the growth rate
1+r is the growth factor
in this problem, the formula becomes:
3000 = p * (1 + .06/4) ^ (4 * 2) which becomes:
3000 = p * 1.015 ^ 8
dolve for p to get:
p = 3000 / (1.015 ^ 8) = 2663.133371.
confirm by replace p with that in the original equation and solving for f to get;
f = 2663.133371 * 1.015 ^ 8 = 3000.
this confirms the value of p is correct.